Calculate the potential for the following redox reaction:
2Fe3+(aq) + Sn2+(aq) ↔ Sn4+(aq) + 2Fe2+(aq)
for a solution in which [Fe3+] = 0.050 M, [Fe2+] = 0.030 M, [Sn2+] = 0.015 M and
[Sn4+] = 0.020 M. Fe3+/ Fe2+ = 0.771 V; Sn4+/ Sn2+ = 0.154 V.
Also show the balanced oxidation and reduction half‐reactions.
1
Expert's answer
2020-06-25T07:46:08-0400

Fe3+ + e- = Fe2+

Sn4+ + 2e- = Sn2+

E = Eo + (RT / zF)ln(ox / red)

EFe3+/ Fe2+ = 0.784 V

ESn4+/ Sn2+ = 0.161 V

E = EFe3+/ Fe2+ - ESn4+/ Sn2+ = 0.784 - 0.161 = 0.623 V.